# The Implication Problem for Functional Dependencies and Variants of Marginal Distribution Equivalences

### v1

@@ 1. Introduction
This paper is an extended version of a conference paper of the same name (Hirvonen, 2022). The axiomatization was introduced in the first version, in which it was also noted that it contains the so-called $k$-cycle rules for all odd natural numbers $k$, which make the axiomatization infinite. In this extended version, we see that this was not caused just by an unfortunate choice of axioms, since we include a proof which shows that there can be no finite axiomatization. This version also contains new results for the subclass without unary marginal distribution equivalences, i.e., the class that contains only functional dependencies and unary marginal identities. We show that for this subclass the implication problem can be simulated with functional dependencies and unary inclusion dependencies. The idea is to simulate each marginal identity $x\approx y$ with two inclusion dependencies $x\subseteq y$ and $y\subseteq x$. Even so, note that there is no semantical equivalence between $x\approx y$ and the corresponding two inclusion dependencies, because two columns $x$ and $y$ may contain exactly the same values but in different quantities. In the earlier version, we constructed a polynomial time algorithm for the full class of functional dependencies and unary marginal identities and marginal distribution equivalences. Since the implication problem for functional dependencies and unary inclusion dependencies is known to be in polynomial time, we obtain from the simulation result a simpler proof for the fact that the implication problem for the subclass is in polynomial time.
@@
A class of dependencies is said to have Armstrong relations if for any finite set $\Sigma$ of dependencies from the class, there is a relation that satisfies a dependency $\sigma$ in the class if and only if $\sigma$ is true for every relation that satisfies $\Sigma$. An Armstrong relation can be used as a convenient representation of a dependency set. If a class of dependecies has Armstrong relations, then the implication problem for a fixed set $\Sigma$ and an arbitrary dependency $\sigma$ is reduced to checking whether the Armstrong relation for $\Sigma$ satisfies $\sigma$. When Armstrong axiomatized functional dependence in (Armstrong, 1974), he also implicitly proved that the class has Armstrong relations (Fagin, 1982). Unfortunately, there are sets of functional dependecies for which the size of a minimal Armstrong relation is exponential in the number of variables (attributes) of the set (Beeri et al., 1984). Armstrong relations can still be useful in practice (Langeveldt and Link, 2010),(Mannila and Räihä, 1986).
@@
Luckily, in our case we can utilize the implication problem for functional dependencies and unary inclusion dependencies which has been axiomatized in (Cosmadakis et al., 1990) for both finite and unrestricted implication. The axiomatization for finite implication is proved to be complete by constructing relations with the help of multigraphs that depict the dependecies between variables. This approach has also been applied to unary functional dependence, unary inclusion dependence, and independence (Hannula and Link, 2018). Our approach is similar, but since we are working in the probabilistic setting, we have to be careful that our construction also works with the two kinds of variants of unary marginal distribution equivalences.
@@ 2. Preliminaries
$|\mathbb{X}_{\bar{x}=\bar{a}}|:=\sum_{\begin{subarray}{c}s(\bar{x})=\bar{a}\\ s\in X\end{subarray}}\mathbb{X}(s).$
@@
Before defining the semantics for the atoms, we need to introduce the notion of a multiset. A multiset is a pair $(B,m)$ where $B$ is a set, and $m\colon B\to\mathbb{N}$ is a multiplicity function. The function $m$ determines for each element $b\in B$ how many multiplicities of $b$ the multiset $(B,m)$ contains. We often denote multisets using double wave brackets, e.g., $(B,m)=\{\{0,1,1\}\}$ when $B=\{0,1\}$ and $m(n)=n+1$ for all $n\in B$.
@@
Example 2.1 ().
Consider the probabilistic team $\mathbb{X}$ depicted in Table 1. The following are some examples of atoms that are (or are not) satisfied in $\mathbb{X}$. Functional dependencies: $\mathbb{X}\models\mathop{=\!}(x,w)$, $\mathbb{X}\models\mathop{=\!}(w)$, and $\mathbb{X}\not\models\mathop{=\!}(x,y)$. Inclusion dependencies: $\mathbb{X}\models x\subseteq y$, $\mathbb{X}\models y\subseteq x$, and $\mathbb{X}\not\models x\subseteq w$. Marginal identities: $\mathbb{X}\models x\approx z$ and $\mathbb{X}\not\models x\approx y$. Marginal distribution equivalences: $\mathbb{X}\models x\approx^{*}y$ and $\mathbb{X}\models x\approx^{*}w$.
@@ 3. Axiomatization for FDs+UMIs+UMDEs
In this section, we present an axiomatization for dependency, unary marginal identity, and unary marginal distribution equivalence atoms. The axiomatization is infinite and it will be shown in section 3.4 that there is no finite axiomatization.
@@ 3.1. Axioms
For the unary functional dependencies and unary marginal distribution equivalencies, we have the cycle rule:
$\text{If }\mathop{=\!}(x_{0},x_{1})\text{ and }x_{1}\approx^{*}x_{2}\text{ and }\dots\text{ and }\mathop{=\!}(x_{k-1},x_{k})\text{ and }x_{k}\approx^{*}x_{0},$
@@
for all $k\in\{1,3,5,\dots\}$. Note that the cycle rule is not a single axiom but an infinite axiom schema, making our axiomatization infinite. The following inference rule is as a useful special case of the cycle rule
@@
Let $\Sigma\cup\{\sigma\}$ be a set of atoms and $\mathcal{A}$ an axiomatization, i.e., a set of axioms. We write $\Sigma\vdash_{\mathcal{A}}\sigma$ iff $\sigma$ can be derived from $\Sigma$ by using the axiomatization $\mathcal{A}$. We denote by $\mathsf{cl}_{\mathcal{A}}(\Sigma)$ the set of atoms obtained by closing $\Sigma$ under the axioms $\mathcal{A}$, i.e., for all $\sigma$ with variables from $D$, $\sigma\in\mathsf{cl}_{\mathcal{A}}(\Sigma)$ iff $\Sigma\vdash_{\mathcal{A}}\sigma$. We say that an axiomatization $\mathcal{A}$ is sound if for any set $\Sigma\cup\{\sigma\}$ of dependencies, $\Sigma\vdash_{\mathcal{A}}\sigma$ implies $\Sigma\models\sigma$. Correspondingly, an axiomatization is complete if for any set $\Sigma\cup\{\sigma\}$ of dependencies, $\Sigma\models\sigma$ implies $\Sigma\vdash_{\mathcal{A}}\sigma$. If the axiomatization $\mathcal{A}$ is clear from the context, we write $\Sigma\vdash\sigma$ and $\mathsf{cl}(\Sigma)$ instead of $\Sigma\vdash_{\mathcal{A}}\sigma$ and $\mathsf{cl}_{\mathcal{A}}(\Sigma)$.
3.2. Soundness of the axioms
In this section, we show that our axiomatization is sound. We only show that the cycle rule is sound; since it is straightforward to check the soundness of other the axioms, we leave out their proofs.
We notice that for all $x,y\in D$, $\mathbb{X}\models\mathop{=\!}(x,y)$ implies $|X(x)|\geq|X(y)|$ and $\mathbb{X}\models x\approx^{*}y$ implies $|X(x)|=|X(y)|$. Suppose now that the precedent of the cycle rule holds for $\mathbb{X}$. It then follows that
@@ 3.3. Completeness and Armstrong relations
In this section, we show that our axiomatization is complete. We do this by showing that for any set $\Sigma$ of FDs, UMIs, and UMDEs, there is a probabilistic team $\mathbb{X}$ such that $\mathbb{X}\models\sigma$ iff $\sigma\in\mathsf{cl}(\Sigma)$. Note that proving this implies completeness: if $\sigma\not\in\mathsf{cl}(\Sigma)$ (i.e. $\Sigma\not\vdash\sigma$), then $\mathbb{X}\not\models\sigma$. Since $\mathbb{X}\models\Sigma$, we have $\Sigma\not\models\sigma$. By doing the proof this way, we obtain Armstrong relations because the constructed team $\mathbb{X}$ is an Armstrong relation [^2] Note that a probabilistic team is actually not a relation but a probability distribution. Therefore, to be exact, instead of Armstrong relations, we should speak of Armstrong models, which is a more general notion introduced in (Fagin, 1982). In our setting, the Armstrong models we construct are uniform distributions over a relation, so each model is determined by a relation, and it suffices to speak of Armstrong relations. for $\Sigma$.
We first define a multigraph in which different-colored edges correspond to different types of dependencies between variables. By using the properties of the graph, we can construct a suitable team, which can then be made into a probabilistic team by taking the uniform distribution over the assignments.
Definition 3.1 ().
@@
    Let $\bar{x}=x_{0}\dots x_{n}$. Suppose that we have $\mathop{=\!}(z,x_{0}\dots x_{k})$ and $\mathop{=\!}(z,x_{k+1})$ for some $10\leq k\leq n-$ 1. By FD1 and FD3, we have $\mathop{=\!}(z,zz)$, $\mathop{=\!}(zz,x_{k+1}z)$, and $\mathop{=\!}(x_{k+1}z,zx_{k+1})$. Then by using FD2 twice, we obtain $\mathop{=\!}(z,zx_{k+1})$. Next, by FD3, we also have $\mathop{=\!}(zx_{k+1},x_{0}\dots x_{k}x_{k+1})$, and then by using FD2, we obtain $\mathop{=\!}(z,x_{0}\dots x_{k+1})$. Since we have $\mathop{=\!}(z,x_{j})$ for all $0\leq j\leq n$, this shows that $\mathop{=\!}(z,\bar{x})$. By using FD2 to $\mathop{=\!}(z,\bar{x})$ and $\mathop{=\!}(\bar{x},y)$, we obtain $\mathop{=\!}(z,y)$.
We next give a unique number to each strongly connected component of $G(\Sigma)$. The numbers are assigned such that the number of a descendant component is always greater that the number of its ancestor component. We call these numbers scc-numbers and denote by $\mathsf{scc}(x)$ the scc-number of the strongly connected component that the vertex $x$ belongs to.
Definition 3.3 ().
Let $\Sigma$ and $G(\Sigma)$ be as in Lemma 3.2 and assign an scc-numbering to $G(\Sigma)$. Let $D$ be the set of the variables appearing in $\Sigma$. We construct a team $X$ over the variables $D$ as follows:
@@
The next lemma shows that team we constructed above already has many of the wanted properties:
Lemma 3.4 ().
Let $\Sigma$, $G(\Sigma)$, and $X$ be as in Definition 3.3. Define a probabilistic team $\mathbb{X}$ as the uniform distribution over $X$. Then the following statements hold:
@@
    Suppose that $\sigma=\mathop{=\!}(\bar{x},\bar{y})\in\Sigma$. We may assume that $\bar{x}$,$\bar{y}$ do not contain variables $z$ for which $\mathop{=\!}(z)\in\Sigma$, as it is easy to show that $\mathbb{X}\models\mathop{=\!}(\bar{x},\bar{y})$ implies $\mathbb{X}\models\mathop{=\!}(\bar{x}z,\bar{y})$ and $\mathbb{X}\models\mathop{=\!}(\bar{x},\bar{y}z)$ for any such variable $z$. For a contradiction suppose that $\mathbb{X}\not\models\sigma$. Then there are tuples $s,s^{\prime}$ that violate $\sigma$. By our assumption, the only repeated number in each column of variables $\bar{x}$,$\bar{y}$ is 0. Thus $s(\bar{x})=s^{\prime}(\bar{x})=\bar{0}$, and either $s(\bar{y})\neq\bar{0}$ or $s^{\prime}(\bar{y})\neq\bar{0}$. Assume that $s(\bar{y})\neq\bar{0}$. Then there is $y_{0}\in\mathsf{var}(\bar{y})$ such that $s(y_{0})\neq 0$. By our construction, tuple $s$ corresponds to a red clique. This means that in $s$, each variable that is 0, is functionally determined by every variable of the corresponding red clique. Since $\Sigma$ is closed under FD rules, there is a variable $z$ such that $s(z)=0$ and $\mathop{=\!}(z,\bar{x})\in\Sigma$. Then also $\mathop{=\!}(z,y_{0})\in\Sigma$, and thus, by the construction, $s(y_{0})=0$, which is a contradiction.

### v2

@@ 1. Introduction
This paper is an extended version of a conference paper of the same title (Hirvonen, 2022). The main results are divided in sections as follows. In Section 3, we introduce the axiomatization and show that it is sound and complete. The axioms contain the so-called $k$-cycle rules for all odd natural numbers $k$, which make the axiomatization infinite. We also demonstrate that this is not due to an inappropriate choice of axioms, since the section includes a proof which shows that there can be no finite axiomatization. Section 4 contains results for the subclass without unary marginal distribution equivalences, i.e., the class that contains only functional dependencies and unary marginal identities. We show that for this subclass the implication problem can be simulated with functional dependencies and unary inclusion dependencies. Since the implication problem for functional dependencies and unary inclusion dependencies is known to be in polynomial time, we obtain from the simulation result a proof for the fact that the implication problem for the subclass is in polynomial time. In Section 5, we construct a polynomial time algorithm for the full class of functional dependencies and unary marginal identities and marginal distribution equivalences.
@@
A class of dependencies is said to have Armstrong relations if for any finite set $\Sigma$ of dependencies from the class, there is a relation that satisfies a dependency $\sigma$ in the class if and only if $\sigma$ is true for every relation that satisfies $\Sigma$. An Armstrong relation can be used as a convenient representation of a dependency set. If a class of dependecies has Armstrong relations, then the implication problem for a fixed set $\Sigma$ and an arbitrary dependency $\sigma$ is reduced to checking whether the Armstrong relation for $\Sigma$ satisfies $\sigma$. When Armstrong axiomatized functional dependence in (Armstrong, 1974), he also implicitly proved that the class has Armstrong relations (Fagin, 1982). However, there are sets of functional dependecies for which the size of a minimal Armstrong relation is exponential in the number of variables (attributes) of the set (Beeri et al., 1984). Armstrong relations can still be useful in practice (Langeveldt and Link, 2010),(Mannila and Räihä, 1986).
@@
In our case, we can utilize the implication problem for functional dependencies and unary inclusion dependencies which has been axiomatized in (Cosmadakis et al., 1990) for both finite and unrestricted implication. The axiomatization for finite implication is proved to be complete by constructing relations with the help of multigraphs that depict the dependecies between variables. This approach has also been applied to unary functional dependence, unary inclusion dependence, and independence (Hannula and Link, 2018). Our approach is similar, but since we are working in the probabilistic setting, our construction must take into account the two kinds of variants of unary marginal distribution equivalences.
@@ 2. Preliminaries
$|\mathbb{X}_{\bar{x}=\bar{a}}|:=\sum_{\begin{subarray}{c}s(\bar{x})=\bar{a}\\ s\in X\end{subarray}}\mathbb{X}(s),$
i.e. $|\mathbb{X}_{\bar{x}=\bar{a}}|$ is the marginal probability of that the variables $\bar{x}$ have the values $\bar{a}$ in the probabilistic team $\mathbb{X}$.
@@
Before defining the semantics for the atoms, we need to introduce the notion of a multiset. A multiset is a pair $(B,m)$ where $B$ is a set, and $m\colon B\to\mathbb{N}$ is a multiplicity function. The function $m$ determines for each element $b\in B$ how many multiplicities of $b$ the multiset $(B,m)$ contains. We often denote multisets using double wave brackets, e.g., $(B,m)=\{\{0,1,1\}\}$ when $B=\{0,1\}$, and $m$ is such that $m(0)=1$ and $m(1)=2$.
@@
Example 2.1.
Consider the probabilistic team $\mathbb{X}$ depicted in Table 1. The following are some examples of atoms that are (or are not) satisfied in $\mathbb{X}$. Functional dependencies: $\mathbb{X}\models\mathop{=\!}(x,w)$ holds because whenever two rows agree on a value of $x$, they also agree on $w$. Also $\mathbb{X}\models\mathop{=\!}(w)$ holds since the value of $w$ is 0 on every row. However, we have $\mathbb{X}\not\models\mathop{=\!}(x,y)$ because on the first two rows the values of $x$ are both 0, but the value of $y$ is 0 on the first row and 1 on the second.
Inclusion dependencies: $\mathbb{X}\models x\subseteq y$ holds because all the values that appear in the $x$ column (i.e. 0 and 1) also appear in the $y$ column. Since the only values in the $y$ column are 0 and 1, we also have $\mathbb{X}\models y\subseteq x$. However $\mathbb{X}\not\models x\subseteq w$ because the value 1 appears in $x$ column, but not in $w$.
Marginal identities: $\mathbb{X}\models x\approx z$ holds because the marginal distributions of $x$ and $z$ are identical: probability of 0 is $2/3$ and the probability of 1 is $1/3$. Since the marginal distribution for $y$ is different: probability of 0 is $1/3$ and the probability of 1 is $2/3$, we have $\mathbb{X}\not\models x\approx y$.
Marginal distribution equivalences: $\mathbb{X}\models x\approx^{*}y$ and $\mathbb{X}\models x\approx^{*}z$ both hold because the multiset of marginal probabilities is the same $\{\{1/3,2/3\}\}$ for all three variables $x$, $y$, and $z$. However, $\mathbb{X}\not\models x\approx^{*}w$ because $\{\{1/3,2/3\}\}\neq\{\{1\}\}$.
Note that the example demonstrates that a marginal identity implies the corresponding marginal distribution equivalence, but not the other way around.
@@ 3. Axiomatization for FDs+UMIs+UMDEs
In this section, we present an axiomatization for the implication problem of dependency, unary marginal identity, and unary marginal distribution equivalence atoms. The axiomatization is infinite and it will be shown in Section 3.3 that there is no finite axiomatization.
@@ 3.1. Axioms
In Armstrong’s original axiomatization, variables represent sets, whereas in this paper, we consider ordered sequences. Note that the difference is not important here because the condition $\mathsf{var}(\bar{y})\subseteq\mathsf{var}(\bar{x})$ in FD1 together with FD2 ensures that the order of variables in a tuple does not matter: any tuple $\bar{x}$ appearing in a functional dependency atom can be replaced by another tuple $\bar{x}^{\prime}$ such that $\mathsf{var}(\bar{x})=\mathsf{var}(\bar{x}^{\prime})$.
@@
Let $k\in\{1,3,5,\dots\}$ and define function $S_{k}\colon\{0,1\dots,k\}\to\{0,1\dots,k\}$ as follows
$S_{k}(i)=\begin{cases}i+1,\text{ when }0\leq i<k\\ 0,\text{ when }i=k.\end{cases}$
For the unary functional dependencies and unary marginal distribution equivalencies, we have the k-cycle rule for all $k\in\{1,3,5,\dots\}$:
If =($\text{If }\mathop{=\!}(x_{i},x_{S_{k}(i)})\text{ for all even }0\leq i\leq k-1\text{ and }x_{j}\approx^{*}x_{S_{k}(j)}\text{ for all odd }1\leq j\leq k,$
@@
The axiom is called a $k$-cycle rule because its antecedent forms a cycle in which FDs and UMDEs alternate. Note that the cycle rules are an infinite axiom schema, making our axiomatization infinite. The following inference rule is as a useful special case of the cycle rule
@@
Let $\Sigma\cup\{\sigma\}$ be a set of atoms and $\mathcal{A}$ an axiomatization, i.e., a set of axioms. Let $D$ be the set of variables that appear in $\Sigma$. We write $\Sigma\vdash_{\mathcal{A}}\sigma$ iff $\sigma$ can be derived from $\Sigma$ by using the axiomatization $\mathcal{A}$. We denote by $\mathsf{cl}_{\mathcal{A}}(\Sigma)$ the set of atoms obtained by closing $\Sigma$ under the axioms $\mathcal{A}$, i.e., for all $\sigma$ with variables from $D$, $\sigma\in\mathsf{cl}_{\mathcal{A}}(\Sigma)$ iff $\Sigma\vdash_{\mathcal{A}}\sigma$. We say that an axiomatization $\mathcal{A}$ is sound if for any set $\Sigma\cup\{\sigma\}$ of dependencies, $\Sigma\vdash_{\mathcal{A}}\sigma$ implies $\Sigma\models\sigma$. Correspondingly, an axiomatization is complete if for any set $\Sigma\cup\{\sigma\}$ of dependencies, $\Sigma\models\sigma$ implies $\Sigma\vdash_{\mathcal{A}}\sigma$. If the axiomatization $\mathcal{A}$ is clear from the context, we write $\Sigma\vdash\sigma$ and $\mathsf{cl}(\Sigma)$ instead of $\Sigma\vdash_{\mathcal{A}}\sigma$ and $\mathsf{cl}_{\mathcal{A}}(\Sigma)$.
The axiomatization defined above is sound. We only show that the cycle rules are sound; since it is straightforward to check the soundness of other the axioms, we leave out their proofs.
We notice that for all $x,y\in D$, $\mathbb{X}\models\mathop{=\!}(x,y)$ implies $|X(x)|\geq|X(y)|$ and $\mathbb{X}\models x\approx^{*}y$ implies $|X(x)|=|X(y)|$. Suppose now that the antecedent of the $k$-cycle rule holds for $\mathbb{X}$. It then follows that
@@
Note that $k$-cycle rules are not in general sound for equivalence relations. Consider the equivalence relation $\approx_{\max}$ such that $\mathbb{X}\models x\approx_{\max}y$ if and only if $\max\{|\mathbb{X}_{x=a}|\mid a\in X(x)\}=\max\{|\mathbb{X}_{y=a}|\mid a\in X(y)\}$. It is easy to see that $\approx_{\max}$ is indeed an equivalence relation. However, the probabilistic team $\mathbb{X}$ in Table 2 demonstrates that $\{\mathop{=\!}(x_{0},x_{1}),x_{1}\approx_{\max}x_{2},\mathop{=\!}(x_{2},x_{3}),x_{3}\approx_{\max}x_{0}\}\not\models\mathop{=\!}(x_{1},x_{0})$, i.e., the $3$-cycle rule for UFDs and $\approx_{\max}$-atoms is not sound. This is because $\max\{|\mathbb{X}_{x_{i}=a}|\mid a\in X(x_{i})\}=1/2$ for all $0\leq i\leq 3$, and therefore both of the atoms $x_{1}\approx_{\max}x_{2}$ and $x_{3}\approx_{\max}x_{0}$ are satisfied. Since it is easy to check that $\mathop{=\!}(x_{0},x_{1})$ and $\mathop{=\!}(x_{2},x_{3})$ are satisfied, but $\mathop{=\!}(x_{1},x_{0})$ is not, we obtain $\mathbb{X}\models\{\mathop{=\!}(x_{0},x_{1}),x_{1}\approx_{\max}x_{2},\mathop{=\!}(x_{2},x_{3}),x_{3}\approx_{\max}x_{0}\}$ and $\mathbb{X}\not\models\mathop{=\!}(x_{1},x_{0})$.
Table 2. The probabilistic team $\mathbb{X}$.
$x_{0}$ $x_{1}$ $x_{2}$ $x_{3}$ $\mathbb{X}$
00 00 00 00 1/4
$1$ 00 00 00 1/4
$2$ $1$ $1$ $1$ 1/4
$2$ $1$ $2$ $1$ 1/4
@@ 3.2. Completeness and Armstrong relations
In this section, we show that our axiomatization is complete and the class of FDs, UMIs and UMDEs has Armstrong relations. Recall that for any finite set $\Sigma$ of dependencies from the class, an Armstrong relation is a probabilistic team $\mathbb{X}$ such that for every $\sigma$ from the class, $\mathbb{X}\models\sigma$ if and only if $\Sigma\models\sigma$.
We prove the completeness by showing that for any set $\Sigma$ of FDs, UMIs, and UMDEs, there is a probabilistic team $\mathbb{X}$ such that $\mathbb{X}\models\sigma$ iff $\sigma\in\mathsf{cl}(\Sigma)$. Note that proving this implies completeness: if $\sigma\not\in\mathsf{cl}(\Sigma)$ (i.e. $\Sigma\not\vdash\sigma$), then $\mathbb{X}\not\models\sigma$. Since $\mathbb{X}\models\Sigma$, we have $\Sigma\not\models\sigma$. By doing the proof this way, we obtain Armstrong relations because by completeness and soundness, $\sigma\in\mathsf{cl}(\Sigma)$ iff $\Sigma\models\sigma$, from which it follows that the constructed team $\mathbb{X}$ is an Armstrong relation for $\Sigma$. Note that a probabilistic team is actually not a relation but a probability distribution. Therefore, to be exact, instead of Armstrong relations, we should speak of Armstrong models, which is a more general notion introduced in (Fagin, 1982). In our setting, the Armstrong models we construct are uniform distributions over a relation, so each model is determined by a relation, and it suffices to speak of Armstrong relations.
We consider the following variant of multigraphs with colored edges:
Definition 3.1.
Let $n\in\mathbb{N}$. A mixed multigraph with colored edges is a tuple $G=(V,E_{1},\dots,E_{n})$, where $V$ is the set of vertices and each $E_{i}$, $1\leq i\leq n$, is either
  • (i)
    a set of ordered pairs $(x,y)\in V\times V$ (directed edge) or
  • (ii)
    a set of subsets $\{x,y\}\subseteq V$ (undirected edge).
The members of each $E_{i}$ are called edges of the color $i$, and each $(V,E_{i})$ is called the $i$-colored subgraph of $G$.
We now define a multigraph in which different-colored edges correspond to different types of dependencies between variables. By using the properties of the graph, we can construct a suitable team, which can then be made into a probabilistic team by taking the uniform distribution over the assignments.
Definition 3.2.
@@
Example 3.3.
Let $\Delta=\{\mathop{=\!}(x_{0},x_{1}),x_{1}\approx x_{2},\mathop{=\!}(x_{2},x_{3}),x_{3}\approx^{*}x_{0},\mathop{=\!}(x_{4}),\mathop{=\!}(x_{5})\}$, and $\Sigma=\mathsf{cl}(\Delta)$. The graph $G(\Sigma)$ constructed as in Definition 3.2 is depicted Figure 1.
$x_{0}$$x_{1}$$x_{2}$$x_{3}$$x_{4}$$x_{5}$
Definition 3.4 (Strongly connected component).
A multigraph $G=(V,E_{1},\dots,E_{n})$ is strongly connected if for all $x,y\in V$ $y$ is reachable from $x$, i.e., there is a sequence of vertices $v_{1},v_{2},\dots,v_{m}\in V$ such that $v_{1}=x$, $v_{m}=y$, and for all $1\leq j\leq m$, $(v_{j},v_{j+1})\in E_{i}$ or $\{v_{j},v_{j+1}\}\in E_{i}$ for some $1\leq i\leq n$. A set of vertices $S\subseteq V$ is said to be a strongly connected component of $G$ if its induced subgraph is a maximal strongly connected subgraph of $G$.
Definition 3.5 (Clique).
Let $G$ a multigraph $(V,E_{1},\dots,E_{n})$ and $i$ be such that $E_{i}$ is a set of undirected edges. An i-colored clique in $G$ is a set of vertices $C\subseteq V$ such that for all $x,y\in C$, $\{x,y\}\in E_{i}$.
@@
    Let $\bar{x}=x_{0}\dots x_{n}$. Each vertex $x_{j}$, $0\leq j\leq n$, has a common ancestor $z$ in the red subgraph of $G(\Sigma)$ and the red subgraph is transitive by part (ii), so $z$ is also a direct ancestor of each $x_{j}$, i.e. $\mathop{=\!}(z,x_{j})\in\Sigma$. Suppose that we have both $\mathop{=\!}(z,x_{0}\dots x_{k})$ and $\mathop{=\!}(z,x_{k+1})$ in $\Sigma$ for some $10\leq k\leq n-$ 1. By FD1 and FD3, we have $\mathop{=\!}(z,zz)$, $\mathop{=\!}(zz,x_{k+1}z)$, and $\mathop{=\!}(x_{k+1}z,zx_{k+1})$ in $\Sigma$. Then by using FD2 twice, we obtain $\mathop{=\!}(z,zx_{k+1})\in\Sigma$. Next, by FD3, we also have $\mathop{=\!}(zx_{k+1},x_{0}\dots x_{k}x_{k+1})\in\Sigma$, and then by using FD2, we obtain $\mathop{=\!}(z,x_{0}\dots x_{k+1})\in\Sigma$. Since we have $\mathop{=\!}(z,x_{j})\in\Sigma$ for all $0\leq j\leq n$, this shows that $\mathop{=\!}(z,\bar{x})\in\Sigma$. By using FD2 to $\mathop{=\!}(z,\bar{x})\in\Sigma$ and $\mathop{=\!}(\bar{x},y)\in\Sigma$, we obtain $\mathop{=\!}(z,y)\in\Sigma$, i.e., there is a red edge from $x$ to $y$.
Example 3.7.
Consider the graph $G(\Sigma)$ of Example 3.3 depicted in Figure 1. The strongly connected components of $G(\Sigma)$ are $\{x_{0},x_{1},x_{2},x_{3}\}$ and $\{x_{4},x_{5}\}$, the maximal red cliques are $\{x_{0},x_{1}\}$, $\{x_{2},x_{3}\}$, and $\{x_{4},x_{5}\}$, the maximal blue cliques are $\{x_{0},x_{1},x_{2},x_{3}\}$ and $\{x_{4},x_{5}\}$ and the maximal black cliques are $\{x_{0}\}$, $\{x_{1},x_{2}\}$, $\{x_{3}\}$, $\{x_{4}\}$, and $\{x_{5}\}$.
Let $S$ and $S^{\prime}$ be two different strongly connected components of a graph. We say that $S^{\prime}$ is a descendant component of $S$ if for every $v\in S$ and $v^{\prime}\in S^{\prime}$, the vertex $v^{\prime}$ is a descendant of $v$. Since this means that for every $v\in S$ and $v^{\prime}\in S^{\prime}$, the vertex $v$ is an ancestor of $v^{\prime}$, we also say that $S$ is an ancestor component of $S^{\prime}$.
We next give a unique number to each strongly connected component of $G(\Sigma)$. The numbers are assigned such that the number of a descendant component is always greater than the number of its ancestor component. We call these numbers scc-numbers and denote by $\mathsf{scc}(x)$ the scc-number of the strongly connected component that the vertex $x$ belongs to. Note that if $G(\Sigma)$ has two strongly connected components without an edge between them, there exist several possible scc-numberings. In this case, we just choose one of the possible numberings as our fixed scc-numbering.
Definition 3.8.
Let $\Sigma$ and $G(\Sigma)$ be as in Lemma 3.6 and assign an scc-numbering to $G(\Sigma)$. Let $D$ be the set of the variables appearing in $\Sigma$. We define a probabilistic team $\mathbb{X}$ as the uniform distribution over $X$, where the team $X$ over the variables $D$ is constructed as follows:
@@
The zero tuple is added in step (i) to make sure that any variable $x$ determines functionally only those variables that appear in the same maximal red clique as $x$ or in some its red descendant cliques. This can be achieved by adding in step (ii), for each strongly connected component, new tuples where zeroes appear in positions according to the maximal red cliques. This is described in more detail in the proof of Lemma 3.10 (i) & (ii).
The purpose of counting in (iib) is to ensure that the columns for variables in the same strongly connected component have the same number of zeroes. Since every pair of variables in the same strongly connected component has an UMDE between them, the same number of zeroes ensures that when we later fill in the empty positions without repeating values, these UMDEs will hold in the team. Note that by this counting, we also ensure that columns that correspond to variables in different strongly connected components have different numbers of zeroes. This ensures that only the inferable UMDEs are satisfied.
When we fill in the empty positions in step (iv), the chosen values are determined by the maximal black cliques, so only the inferable UMIs hold. Note that if there are constant variables, their values will also be renamed in step (v) for the same reason.
Example 3.9.
Let $\Sigma$ be as in Example 3.7 and the probabilistic team $\mathbb{X}$ as in Definition 3.8. The team $\mathbb{X}$ is depicted in Table 3.
Table 3. The probabilistic team $\mathbb{X}$ of Example 3.9 constructed as in Definition 3.8.
$x_{0}$ $x_{1}$ $x_{2}$ $x_{3}$ $x_{4}$ $x_{5}$ $\mathbb{X}$
00 00 00 00 00 $1$ 1/4
00 00 $1$ $1$ 00 $1$ 1/4
$1$ $1$ 00 00 00 $1$ 1/4
$2$ $3$ $3$ $4$ 00 $1$ 1/4
The next lemma shows that the team we constructed above already has many of the wanted properties:
Lemma 3.10.
Let $\Sigma$, $G(\Sigma)$, and $\mathbb{X}$ be as in Definition 3.8. Then the following statements hold:
@@
    Suppose that $\sigma=\mathop{=\!}(\bar{x},\bar{y})\in\Sigma$. We may assume that $\bar{x}$,$\bar{y}$ do not contain variables $z$ for which $\mathop{=\!}(z)\in\Sigma$. This is because for any such variable $z$, $\mathbb{X}\models\mathop{=\!}(\bar{x},\bar{y})$ implies $\mathbb{X}\models\mathop{=\!}(\bar{x}z,\bar{y})$ and $\mathbb{X}\models\mathop{=\!}(\bar{x},\bar{y}z)$. By FD1, we have $\mathop{=\!}(\bar{x}z,\bar{x})$, which together with $\mathop{=\!}(\bar{x},\bar{y})$ implies $\mathop{=\!}(\bar{x}z,\bar{y})$ by FD2. From $\mathop{=\!}(z)$, we obtain $\mathop{=\!}(\bar{y},\bar{y}z)$ by FD3. Then by applying FD2 to $\mathop{=\!}(\bar{x},\bar{y})$ and $\mathop{=\!}(\bar{y},\bar{y}z)$, we obtain $\mathop{=\!}(\bar{x},\bar{y}z)$.
    For a contradiction suppose that $\mathbb{X}\not\models\sigma$. Then there are tuples $s,s^{\prime}$ that violate $\sigma$. By our assumption, the only repeated number in each column of variables $\bar{x}$,$\bar{y}$ is 0. Thus $s(\bar{x})=s^{\prime}(\bar{x})=\bar{0}$, and either $s(\bar{y})\neq\bar{0}$ or $s^{\prime}(\bar{y})\neq\bar{0}$. Assume that $s(\bar{y})\neq\bar{0}$. Then there is $y_{0}\in\mathsf{var}(\bar{y})$ such that $s(y_{0})\neq 0$. By our construction, tuple $s$ corresponds to a red clique. This means that in $s$, each variable that is 0, is functionally determined by every variable of the corresponding red clique. Hence, there is a variable $z$ such that $s(z)=0$ and $z$ is a common ancestor of vertices $\bar{x}$ in the red subgraph. As $\mathop{=\!}(\bar{x},\bar{y})\in\Sigma$ and, by FD1, $\mathop{=\!}(\bar{y},y_{0})\in\Sigma$, we obtain $\mathop{=\!}(\bar{x},y_{0})\in\Sigma$ by using FD2. Then by Lemma 3.6 (v), we have $\mathop{=\!}(z,y_{0})\in\Sigma$, and thus, by the construction, $s(y_{0})=0$, which is a contradiction.